# Goldfish: Monolingual Language Models for 350 Languages

### v1

@@ Appendix A Appendix
A.1 Training Dataset Details
@@
As described in §3.1, we merge the text datasets compiled in Chang et al. (2023), Glot500 (Imani et al., 2023), and MADLAD-400 (clean split; Kudugunta et al., 2023). These datasets include popular multilingual corpora such as OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021), Wikipedia (Wikipedia, 2024), No Language Left Behind (Costa-jussà et al., 2022), and others. Together, these datasets take advantage of both automatically crawled datasets with automated language identification and targeted datasets manually annotated for specific low-resource languages. All included datasets are publicly available; see Limitations for licensing concerns. Comprehensively, the Goldfish dataset includes:
  • Chang et al. (2023):
@@
    AI4Bharat (AI4Bharat, 2023), AI FOR THAI LotusCorpus (AI FOR THAI, 2023), Arabic Dialects Dataset (El-Haj et al., 2018), AfriBERTa (Ogueji et al., 2021), AfroMAFT (Adelani et al., 2022; Xue et al., 2021), Anuvaad (Anuvaad, 2023), AraBench (Sajjad et al., 2020), Autshumato (Autshumato, 2023) Bloom Library (Leong et al., 2022), CC100 (Conneau et al., 2020), CCNet (Wenzek et al., 2020), CMU Haitian Creole (CMU, 2010), SADiLaR NCHLT corpus (SADiLaR, 2023b), Clarin (Clarin, 2023), DART (Alsarsour et al., 2018), Earthlings (Dunn, 2020), FFR Dataset (FFR Dataset, 2023), GiossaMedia (Góngora et al., 2022, 2021), Glosses (Camacho-Collados et al., 2016), Habibi (El-Haj, 2020), HinDialect (Bafna, 2022), HornMT (HornMT, 2023), IITB (Kunchukuttan et al., 2018), IndicNLP (Nakazawa et al., 2021), Indiccorp (Kakwani et al., 2020), isiZulu (Mabuya et al., 2023), JParaCrawl (Morishita et al., 2020), kinyarwandaSMT (Niyongabo, 2023), LeipzigData (Goldhahn et al., 2012), LINDAT (LINDAT, 2023), Lingala Song Lyrics (Lingala Songs, 2023), LyricsTranslate (LyricsTranslate, 2023), mC4 (Raffel et al., 2020), MTData (Gowda et al., 2021), MaCoCu (Bañón et al., 2022), Makerere MT Corpus (Mukiibi et al., 2022), Masakhane Community (Masakhane, 2023), Mburisano Covid Corpus (SADiLaR, 2023a), Menyo20K (Adelani et al., 2021a), Minangkabau corpora (Koto and Koto, 2020), MoT (Palen-Michel et al., 2022), NLLB seed (Costa-jussà et al., 2022), Nart Abkhaz text (Nart, 2023), OPUS (Tiedemann, 2012), OSCAR (Ortiz Suárez et al., 2019), ParaCrawl (Bañón et al., 2020), Parallel Corpora for Ethiopian Languages (Teferra Abate et al., 2018), Phontron (Neubig, 2011), QADI (Abdelali et al., 2021), Quechua-IIC (Zevallos et al., 2022), SLI GalWeb.1.0 (Agerri et al., 2018), Shami (Abu Kwaik et al., 2018), Stanford NLP (Stanford, 2023), StatMT (Koehn, 2023), TICO (Anastasopoulos et al., 2020), TIL (Mirzakhalov et al., 2021), Tatoeba (Tiedemann, 2020), TeDDi (Moran et al., 2022), Tilde (Rozis and Skadiņš, 2017), W2C (Majliš, 2011), WAT (Nakazawa et al., 2022), WikiMatrix (Schwenk et al., 2021), Wikipedia (Wikipedia, 2024), Workshop on NER for South and South East Asian Languages (Singh, 2008), and XLSum (Hasan et al., 2021).
@@
We start with the corpus from Chang et al. (2023). We then merge the dataset per language with Glot500 for languages that have not yet reached our 1GB maximum (after byte premium scaling). Then, we merge the dataset with MADLAD-400 for languages that have still not reached our 1GB maximum. We also add MADLAD-400 for languages with short average line lengths (less than 25.0 tokens), to make use of MADLAD-400’s longer contiguous sequences. To allow comparisons on popular low-resource language evaluations, we exclude FLORES-200 (Costa-jussà et al., 2022) and AmericasNLI (Ebrahimi et al., 2022) from all dataset merging. For each dataset, we exclude languages that contain only Bible data. Because there is likely significant overlap between different dataset sources, we deduplicate repeated sequences of 100 UTF-8 bytes for each language (Lee et al., 2022).
@@
Additionally, we drop several redundant language codes:
@@
The resulting 350 Goldfish languages cover five continents, 28 top-level language families (Hammarström et al., 2023), and 32 scripts (writing systems). All languages for which Goldfish models are available are listed in Table LABEL:tab:lang-list. We include the language name, ISO 639-3 language code, ISO 15924 script code, estimated byte premium, dataset size after byte premium scaling, dataset size in tokens, and proportion of the dataset from each of our four largest sources. Raw dataset sizes before byte premium scaling can be obtained by multiplying the dataset size after byte premium scaling by the estimated byte premium. Source dataset proportions are reported before deduplication. The reported dataset sizes reflect the dataset for the Goldfish model trained on the maximum amount of data for that language (the 1GB-dataset Goldfish when available, otherwise the full-dataset Goldfish). Reported token counts use the tokenizer for the largest Goldfish model for that language. All dataset statistics can be downloaded at https://github.com/tylerachang/goldfish.
A.2 Pre-Training Details
As described in §3.2, we train monolingual language models for five dataset sizes when available after byte premium scaling: 5MB, 10MB, 100MB, 1GB, and full. The full dataset size (including all available data) is only included if a 1GB dataset is not available for a language. In total, the Goldfish include 350 5MB-dataset models, 288 10MB-dataset models, 166 100MB-dataset models, 83 1GB-dataset models, and 267 full-dataset models (1154 models total). Full hyperparameters are reported in Table 4.
Hyperparameter 5MB,10MB 100MB,1GB,full
@@ A.2 Pre-Training Details
Warmup steps 10% of pre-training
@@
Table 4: Pre-training hyperparameters for Goldfish trained on different dataset sizes (Devlin et al., 2019; Turc et al., 2019; Radford et al., 2018).
Tokenizers.
All tokenizers are trained with vocabulary size 50K (Liu et al., 2019) on the same dataset size as their corresponding model (including byte premium scaling). We use SentencePiece tokenizers (Kudo and Richardson, 2018) with training text randomly sampled from the dataset for the desired language. To avoid memory errors, we limit tokenizer training text to 100MB after byte premium scaling. After tokenizer training, we tokenize each training dataset, concatenating text lines such that each sequence contains exactly 512 tokens. We run tokenization before shuffling and sampling to the desired dataset sizes, so our sequences of 512 tokens preserve contiguous text where possible, although several of our source corpora only exist in shuffled form. Finally, we sample our tokenized datasets to 5MB, 10MB, 100MB, and 1GB after byte premium scaling. [^6] When de-tokenized, the tokenized datasets result in slightly smaller datasets than the original text datasets, because the tokenizer truncates lines to create 512-token sequences. All reported dataset sizes account for this truncation.
@@
All of our models use the GPT-2 architecture (Radford et al., 2019), changing only the number of layers, attention heads, and embedding sizes as in Turc et al. (2019). For the 100MB-, 1GB-, and full-dataset models, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018) (similar to BERT-base and RoBERTa; Devlin et al., 2019; Liu et al., 2019). Because smaller models perform similarly to larger models in low-resource scenarios (Chang et al., 2023), we use the small model size (39M parameters) from Turc et al. (2019) for the 10MB and 5MB dataset sizes.
@@
Language models are pre-trained using the Hugging Face Transformers library (Wolf et al., 2020) and code from Chang and Bergen (2022). We refrain from extensive hyperparameter tuning to avoid biasing our hyperparameters towards English (or any other selected tuning language). Instead, we adopt hyperparameters from previous work with minimal modifications. To match the setup of our models and to prevent overfitting, we select hyperparameters based on models with fairly small training datasets relative to modern standards. Specifically, following BERT (Devlin et al., 2019), we use learning rate 1e-4 for the 125M-parameter models (the same as RoBERTa for small batch sizes; Liu et al., 2019; GPT-1 uses learning rate 2.5e-4; Radford et al., 2018). Based on initial results using randomly-sampled languages, we find that learning rate 1e-4 also works well for the 39M-parameter models; this is in line with Chang et al. (2023), who find that learning rate 2e-4 works well for small models, and smaller learning rates reduce the speed of any potential overfitting.
We train each model for 10 epochs of the training data; multiple epochs of pre-training is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but pre-training on more than 10 epochs often leads to overfitting (increases in eval loss) in the 5MB scenarios. For batch sizes, following GPT-1 (most similar to our models; Radford et al., 2018), we use batch size 64 (64$\times$512 = 32K tokens) for the 1GB-dataset models. We find that these larger batch sizes lead to overfitting for small datasets, so we use batch sizes 4, 8, and 32 for 5MB-, 10MB-, and 100MB-dataset models respectively (determined based on initial experiments with randomly-sampled languages). These correspond to batches of 2K, 4K, or 16K tokens. For full-dataset models, we use the batch size that would be used if rounding the dataset size down to 5MB, 10MB, or 100MB (recall that we do not train a full-dataset model when the 1GB dataset is available for a language).
Compute costs.
All language model pre-training runs together take a total of $1.65\times 10^{20}$ FLOPs. This is less than 1/1900×1/1900$\times$1 / 1900 × the computation used to train the original 175B-parameter GPT-3 model (Brown et al., 2020; $3.14\times 10^{23}$ FLOPs). Models are each trained on one NVIDIA GeForce GTX TITAN X, GeForce RTX 2080 Ti, TITAN Xp, Quadro P6000, RTX A4500, RTX A5000, or RTX A6000 GPU. In total, Goldfish pre-training takes the equivalent of approximately 15600 A6000 GPU hours. Inference for FLORES perplexities and reasoning benchmarks takes approximately 250 A6000 GPU hours (primarily due to the large multilingual models used for comparison). Dataset merging, deduplication, and tokenization takes approximately 1600 CPU core hours.
A.3 FLORES Evaluation Details
@@
FLORES log-perplexities for all models and languages are reported in Table LABEL:tab:ppl-table. For Goldfish models, we report the log-perplexity for the model trained on the largest dataset for the language (i.e. the 1GB-dataset model when available, otherwise the full-dataset model). Log-perplexities of the 5MB-, 10MB-, 100MB-, and 1GB-dataset models specifically are available at https://github.com/tylerachang/goldfish.
Bigram model details.
For each FLORES language, we train a bigram model on the entire Goldfish dataset for that language, up to 1GB after byte premium scaling §3.1. The bigram model computes the probability of each token $w_{i}$ as $P(w_{i}|w_{i-1})$, computed based on raw bigram counts in the tokenized Goldfish dataset. The tokenizer is the same as the Goldfish tokenizer for that dataset (i.e. the 1GB-dataset model when available, or the full-dataset model). When a bigram is not observed in the dataset, we use backoff to unigram probability with a penalty multiplier of $\lambda=0.40$ (i.e. “stupid backoff”; Brants et al., 2007). We do not consider $n$-grams for $n>2$ because those $n$-grams often resort to backoff and are therefore much more sensitive to the backoff penalty term $\lambda$.

### v2

@@ Appendix A Appendix
A.1.  Training Dataset Details
@@
As described in §3.1, we merge the text datasets compiled in Chang et al. (2024a), Glot500 (Imani et al., 2023), and MADLAD-400 (clean split; Kudugunta et al., 2023). These datasets include popular multilingual corpora such as OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021), Wikipedia (Wikipedia, 2024), No Language Left Behind (Costa-jussà et al., 2022), and others. Together, these datasets take advantage of both automatically crawled datasets with automated language identification and targeted datasets manually annotated for specific low-resource languages. All included datasets are publicly available; see Limitations for licensing concerns. Comprehensively, the Goldfish dataset includes:
  • Chang et al. (2024a):
@@
    AI4Bharat (AI4Bharat, 2023), AI FOR THAI LotusCorpus (AI FOR THAI, 2023), Arabic Dialects Dataset (El-Haj et al., 2018), AfriBERTa (Ogueji et al., 2021), AfroMAFT (Adelani et al., 2022; Xue et al., 2021), Anuvaad (Anuvaad, 2023), AraBench (Sajjad et al., 2020), Autshumato (Autshumato, 2023) Bloom Library (Leong et al., 2022), CC100 (Conneau et al., 2020b), CCNet (Wenzek et al., 2020), CMU Haitian Creole (CMU, 2010), SADiLaR NCHLT corpus (SADiLaR, 2023b), Clarin (Clarin, 2023), DART (Alsarsour et al., 2018), Earthlings (Dunn, 2020), FFR Dataset (FFR Dataset, 2023), GiossaMedia (Góngora et al., 2022, 2021), Glosses (Camacho-Collados et al., 2016), Habibi (El-Haj, 2020), HinDialect (Bafna, 2022), HornMT (HornMT, 2023), IITB (Kunchukuttan et al., 2018), IndicNLP (Nakazawa et al., 2021), Indiccorp (Kakwani et al., 2020), isiZulu (Mabuya et al., 2023), JParaCrawl (Morishita et al., 2020), kinyarwandaSMT (Niyongabo, 2023), LeipzigData (Goldhahn et al., 2012), LINDAT (LINDAT, 2023), Lingala Song Lyrics (Lingala Songs, 2023), LyricsTranslate (LyricsTranslate, 2023), mC4 (Raffel et al., 2020), MTData (Gowda et al., 2021), MaCoCu (Bañón et al., 2022), Makerere MT Corpus (Mukiibi et al., 2022), Masakhane Community (Masakhane, 2023), Mburisano Covid Corpus (SADiLaR, 2023a), Menyo20K (Adelani et al., 2021a), Minangkabau corpora (Koto and Koto, 2020), MoT (Palen-Michel et al., 2022), NLLB seed (Costa-jussà et al., 2022), Nart Abkhaz text (Nart, 2023), OPUS (Tiedemann, 2012), OSCAR (Ortiz Suárez et al., 2019), ParaCrawl (Bañón et al., 2020), Parallel Corpora for Ethiopian Languages (Teferra Abate et al., 2018), Phontron (Neubig, 2011), QADI (Abdelali et al., 2021), Quechua-IIC (Zevallos et al., 2022), SLI GalWeb.1.0 (Agerri et al., 2018), Shami (Abu Kwaik et al., 2018), Stanford NLP (Stanford, 2023), StatMT (Koehn, 2023), TICO (Anastasopoulos et al., 2020), TIL (Mirzakhalov et al., 2021), Tatoeba (Tiedemann, 2020), TeDDi (Moran et al., 2022), Tilde (Rozis and Skadiņš, 2017), W2C (Majliš, 2011), WAT (Nakazawa et al., 2022), WikiMatrix (Schwenk et al., 2021), Wikipedia (Wikipedia, 2024), Workshop on NER for South and South East Asian Languages (Singh, 2008a), and XLSum (Hasan et al., 2021).
@@
We start with the corpus from Chang et al. (2024a). We then merge the dataset per language with Glot500 for languages that have not yet reached our 1GB maximum (after byte premium scaling). Then, we merge the dataset with MADLAD-400 for languages that have still not reached our 1GB maximum. We also add MADLAD-400 for languages with short average line lengths (less than 25.0 tokens), to make use of MADLAD-400’s longer contiguous sequences. To allow comparisons on popular low-resource language evaluations, we exclude FLORES-200 (Costa-jussà et al., 2022) and AmericasNLI (Ebrahimi et al., 2022) from all dataset merging. For each dataset, we exclude languages that contain only Bible data. Because there is likely significant overlap between different dataset sources, we deduplicate repeated sequences of 100 UTF-8 bytes for each language (Lee et al., 2022).
@@
We also drop several redundant language codes:
@@
The resulting 350 Goldfish languages cover five continents, 28 top-level language families (Hammarström et al., 2023), and 32 scripts (writing systems). All languages for which Goldfish models are available are listed in Table LABEL:tab:lang-list. We include the language name, ISO 639-3 language code, ISO 15924 script code, estimated byte premium, dataset size after byte premium scaling, dataset size in tokens, and proportion of the dataset from each of our four largest sources. Raw dataset sizes before byte premium scaling can be obtained by multiplying the dataset size after byte premium scaling by the estimated byte premium. Source dataset proportions are reported before deduplication. The reported dataset sizes reflect the dataset for the Goldfish model trained on the maximum amount of data for that language (the 1GB-dataset Goldfish when available, otherwise the full-dataset Goldfish). Reported token counts use the tokenizer for the largest Goldfish model for that language. Datasets can be downloaded at: https://huggingface.co/datasets/goldfish-models/fish-food.
A.2.  Contamination Analysis
While we aim to exclude the FLORES dataset from our training datasets, there is still the possibility that FLORES could be contaminated inadvertently due to the presence of web data. To check for any potential contamination of FLORES in the datasets, we tokenize each FLORES sequence and the entire training dataset for each language that is in both Goldfish and FLORES (204 languages). We compute the total number of times that the first 10 tokens of any FLORES example appears in the training dataset for the language. For 72% of languages, no FLORES examples appear in the training dataset at all. For 98% of languages, less than 10 FLORES examples appear in the training dataset (out of over 2000 FLORES examples total). The only two languages with notable FLORES contamination are smo_latn (Samoan; 7155 occurrences of FLORES examples in the training dataset) and knc_arab (Central Kanuri; 371 occurrences of FLORES examples in the training dataset).
A.3.  Pretraining Details
Hyperparameter 5MB, 10MB full, 100MB+
@@ A.3.  Pretraining Details
Warmup steps 10% of pretraining
@@
Table 4: Pretraining hyperparameters for Goldfish trained on different dataset sizes (Devlin et al., 2019; Turc et al., 2019; Radford et al., 2018).
@@
All of our models use the GPT-2 architecture (Radford et al., 2019), changing only the number of layers, attention heads, and embedding sizes as in Turc et al. (2019). For the 100MB-, 1GB-, and full-dataset models, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018) (similar to BERT-base and RoBERTa; Devlin et al., 2019; Liu et al., 2019). Because smaller models perform similarly to larger models in low-resource scenarios (Chang et al., 2024a), we use the small model size (39M parameters) from Turc et al. (2019) for the 10MB and 5MB dataset sizes.
@@
Language models are pretrained using the Hugging Face Transformers library (Wolf et al., 2020) and code from Chang and Bergen (2022). We refrain from extensive hyperparameter tuning to avoid biasing our hyperparameters towards English (or any other selected tuning language). Instead, we adopt hyperparameters from previous work with minimal modifications. To match the setup of our models and to prevent overfitting, we select hyperparameters based on models with fairly small training datasets relative to modern standards. Specifically, following BERT (Devlin et al., 2019), we use learning rate 1e-4 for the 125M-parameter models (the same as RoBERTa for small batch sizes; Liu et al., 2019; GPT-1 uses learning rate 2.5e-4; Radford et al., 2018). Based on initial results using randomly-sampled languages, we find that learning rate 1e-4 also works well for the 39M-parameter models; this is in line with Chang et al. (2024a), who find that learning rate 2e-4 works well for small models, and smaller learning rates reduce the speed of any potential overfitting.
@@
Models are each trained on one NVIDIA GeForce GTX TITAN X, GeForce RTX 2080 Ti, TITAN Xp, Quadro P6000, RTX A4500, RTX A5000, or RTX A6000 GPU. In total, Goldfish pretraining takes the equivalent of approximately 15600 A6000 GPU hours. Inference for FLORES perplexities and reasoning benchmarks takes approximately 250 A6000 GPU hours (primarily due to the large multilingual models used for comparison). Dataset merging, deduplication, and tokenization takes approximately 1600 CPU core hours.
A.4.  FLORES Evaluation Details
In §4.1, we evaluate the Goldfish models, XGLM 4.5B, XGLM 7.5B, BLOOM 7.1B, MaLA-500 10B, and bigram models on FLORES log-perplexity (negative log-likelihood). To avoid confounds from different tokenizers across models, we compute log-perplexities at the sequence level. For fair comparison with multilingual models that need to determine the input language during the early parts of a sequence, we compute log-perplexity of the second half $s_{1}$ of each sequence given the first half $s_{0}$. We then compute the mean over sequences:
In detail, the first and second half of each sequence are determined based on number of characters, so the halfway split is the same for all models considered. We round to the nearest token when the halfway split is in the middle of a subword token. Each model $\mathcal{M}$ then assigns some probability $P_{\mathcal{M}}(s_{1}|s_{0})$ regardless of tokenization, except for rounding the halfway point to the nearest token. The probability for any [UNK] (unknown) token is set to random chance $1/v$ where $v$ is the tokenizer vocabulary size. [^5] Otherwise, for unseen writing systems (e.g. Tibetan script tibt in XGLM), the probability $P(\textrm{[UNK]}|\textrm{[UNK] [UNK] ...})$ is very high, resulting in artificially low perplexities. Setting the [UNK] token probabilities to random chance has very little effect on log-perplexity scores except for the scenario of an unseen writing system. As our final log-perplexity score, we compute the mean negative-log-probability over all FLORES sequences in the target language. Because perplexities generally use geometric means, we use arithmetic means for log-perplexities. The final equation is presented in Equation 2.
The mean FLORES perplexities for each model are reported in Table 1. For Goldfish models, we report the perplexity for the model trained on the largest dataset for the language (i.e. the 1GB-dataset model when available, otherwise the full-dataset model). Perplexities per language for the 5MB-, 10MB-, 100MB-, and 1GB-dataset models specifically are available at https://github.com/tylerachang/goldfish.